# Critical Quantum Sensing: a tutorial on parameter estimation near quantum phase transitions

### v1

@@ XI Conclusions
Critical quantum metrology sits at a frontier where several of the most vibrant areas of modern quantum physics converge. It draws simultaneously on parameter estimation theory, information geometry, and the physics of critical phenomena, providing a common language in which concepts from quantum sensing and many-body theory meet. Within this unifying framework, the universality that emerges at critical points—driven by long-range correlations and diverging susceptibilities—-naturally manifests as an amplification of parameter sensitivity, making criticality an intrinsic resource for metrology. We begin by introducing the fundamentals of parameter estimation theory, and use Ramsey interferometry as a benchmark against which to contrast critical metrology. From there, we build up from simple toy systems such as the harmonic oscillator and Landau–Zener model to paradigmatic many-body systems including the transverse-field Ising and Lipkin–Meshkov–Glick models: illustrating how the quantum state geometry, universal scaling laws, finite-size effects, and dynamical protocols shape the attainable precision. Along the way, we explore the subtleties of multiparameter estimation in critical regimes, as well as the role of finite temperature, before finally connecting these insights to realistic platforms such as driven–dissipative resonators and the ultrastrong light–matter coupling regime, where critical metrology can be tested under experimentally relevant conditions.
@@
XII Outlook
The main challenge in the experimental implementation of critical quantum sensing protocols is to operate systems in proximity of critical points, where the high susceptibility requires high stability of control parameters. Depending on the platform considered, the long timescales can also represent a technical challenge, even if formally the critical slowing down does not prevent achieving the optimal scaling also with respect to time. First experimental implementations already prove the practical feasibility of this approach. For example, many-body critical enhanced metrology for the sensing of external microwave electric fields has been demonstrated in a non-equilibrium Rydberg atomic gas [46]. An adiabatic scheme on a perturbed Ising spin model with a first-order QPT has been implemented with nuclear magnetic resonance techniques [276]. A critical protocol for global quantum sensing has been implemented with a two-qubit superconducting device [277]. A recent experiment demonstrated that quantum criticality in a non-Hermitian topological quantum walk can boost parameter estimation sensitivity, even during the transient dynamics [278]. Finally, a quadratic scaling of the estimation precision with respect to the effective system size has been demonstrated exploiting a driven-dissipative phase transition in a parametric superconducting resonator [65]. On a more general basis, a recent review collates the various methodologies by which many-body systems can be levered for enhanced parameter sensitivity more generally [67]
An emerging theme is that critical metrology is not confined to equilibrium physics. Phenomena such as discrete time crystals [279, 280, 281], illustrate how dynamical phases of matter can also be exploited as metrological resources. This perspective extends the scope of critical metrology well beyond traditional settings. In summary, critical metrology challenges us to move beyond simple paradigms: its success will depend on protocols that combine theoretical elegance with experimental practicality, translating abstract ideas about scaling into real, resource-aware advances in sensing. Looking ahead, several promising research avenues may define the next phase of critical metrology:
@@ XII Outlook
  • Engineering gapped critical-like states. A long-standing challenge in critical metrology is the closing of the energy gap, which slows down adiabatic state preparation. An exciting direction is to engineer nonclassical, entangled eigenstates that mimic the advantageous sensitivity of states near criticality, but retain or even enlarge the energy gap [216]. Such states would combine enhanced precision with fast preparation times, potentially making critical (or perhaps anti-critical) metrology experimentally much more viable.
@@
  • Quantum transport. Experiments with nanoscale devices are capable of probing charge, spin, and heat flow with high precision [283, 284]—where observables such as conductance and current may be highly sensitive to control fields and temperature [285, 286, 287, 288, 289]. These platforms, long used to study many-body phenomena such as Coulomb blockade [290, 291], Kondo physics [283, 292, 293, 294, 295, 197, 296, 297, 298, 299, 300], and quantum Hall effects [301, 302, 303, 304, 305], may act as natural resources for metrology. Furthermore, quantum criticality has already been observed in such systems [306, 307, 308, 309, 310, 311]. Unlike optical schemes [312, 313, 314, 315, 316, 43], transport devices rely on current-based measurements typically operating in strongly interacting, non-Markovian regimes. This necessitates novel approaches to bound ultimate precision via the quantum Fisher information and to quantify precision of experimentally accessible measurements. Recent progress has laid this foundation [56, 317], positioning transport as a natural arena for quantum metrology: being experimentally ubiquitous and supported by mature measurement infrastructure. An exciting avenue for exploration lies in the emerging connection between quantum thermodynamics [318] and metrology, where recent investigations have begun to reveal deep conceptual and practical links [319, 320, 321, 322, 323].
In essence, critical metrology is still in its formative stage. By uniting ideas from equilibrium and non-equilibrium critical phenomena, open-system dynamics, and engineered many-body states, the field has the potential to push quantum sensing far beyond the limits of conventional approaches.
@@ Acknowledgements.
S.F. acknowledges financial support from the foundation Compagnia di San Paolo, grant vEIcolo no. 121319, and from PNRR MUR project PE0000023-NQSTI financed by the European Union – Next Generation EU. U.A. and R.D. acknowledge financial support from the Academy of Finland, grants no. 353832 and 349199. G.M. acknowledges fruitful exchanges with Scott. K.G. would like to acknowledge discussions with Linda and Lidia.

### v2

@@ XII Conclusions
Critical quantum metrology sits at a frontier where several of the most vibrant areas of modern quantum physics converge. It draws simultaneously on parameter estimation theory, information geometry, and the physics of critical phenomena, providing a common language in which concepts from quantum sensing and many-body theory meet. Within this unifying framework, the universality that emerges at critical points—driven by long-range correlations and diverging susceptibilities—-naturally manifests as an amplification of parameter sensitivity, making criticality an intrinsic resource for metrology. We begin by introducing the fundamentals of parameter estimation theory, and use Ramsey interferometry as a benchmark against which to contrast critical metrology. From there, we build up from simple toy systems such as the harmonic oscillator and Landau–Zener model to paradigmatic many-body systems including the transverse-field Ising and Lipkin–Meshkov–Glick models: illustrating how the quantum state geometry, universal scaling laws, finite-size effects, and dynamical protocols shape the attainable precision. Along the way, we explore the subtleties of multiparameter estimation in critical regimes, as well as the role of finite temperature, before finally connecting these insights to realistic platforms such as driven–dissipative resonators, ultrastrong light–matter coupling regime, and strongly correlated electron systems, where critical metrology can be tested under experimentally relevant conditions.
@@
XIII Outlook
The main challenge in the experimental implementation of critical quantum sensing protocols is to operate systems in proximity of critical points, where the high susceptibility requires high stability of control parameters. Depending on the platform considered, the long timescales can also represent a technical challenge, even if formally the critical slowing down does not prevent achieving the optimal scaling also with respect to time. First experimental implementations already prove the practical feasibility of this approach. For example, many-body critical enhanced metrology for the sensing of external microwave electric fields has been demonstrated in a non-equilibrium Rydberg atomic gas [50]. An adiabatic scheme on a perturbed Ising spin model with a first-order QPT has been implemented with nuclear magnetic resonance techniques [304]. A critical protocol for global quantum sensing has been implemented with a two-qubit superconducting device [305]. A 9-qubit superconducting device has been used to test a critical sensing protocol on a Stark-Wannier localization transition [306]. A recent experiment demonstrated that quantum criticality in a non-Hermitian topological quantum walk can boost parameter estimation sensitivity, even during the transient dynamics [307]. Finally, a quadratic scaling of the estimation precision with respect to the effective system size has been demonstrated exploiting a driven-dissipative phase transition in a parametric superconducting resonator [68]. On a more general basis, a recent review collates the various methodologies by which many-body systems can be levered for enhanced parameter sensitivity more generally [69]
An emerging theme is that critical metrology is not confined to equilibrium physics. Phenomena such as discrete time crystals [308, 309, 310, 311] illustrate how dynamical phases of matter can also be exploited as metrological resources. This perspective extends the scope of critical metrology well beyond traditional settings. In summary, critical metrology challenges us to move beyond simple paradigms: its success will depend on protocols that combine theoretical elegance with experimental practicality, translating abstract ideas about scaling into real, resource-aware advances in sensing. Looking ahead, several promising research avenues may define the next phase of critical metrology:
@@ XIII Outlook
  • Engineering gapped critical-like states. A long-standing challenge in critical metrology is the closing of the energy gap, which slows down adiabatic state preparation. An exciting direction is to engineer nonclassical, entangled eigenstates that mimic the advantageous sensitivity of states near criticality, but retain or even enlarge the energy gap [312, 230]. Such states would combine enhanced precision with fast preparation times, potentially making critical (or perhaps anti-critical) metrology experimentally much more viable.
@@
  • Quantum transport. Experiments with nanoscale devices are capable of probing charge, spin, and heat flow with high precision [294, 314]—where observables such as conductance and current may be highly sensitive to control fields and temperature [315, 316, 317, 318, 319]. These platforms, long used to study many-body phenomena such as Coulomb blockade [320, 321], Kondo physics [294, 322, 296, 323, 324, 210, 325, 293, 326, 327, 328], and quantum Hall effects [329, 330, 331, 332, 333], may act as natural resources for metrology. Furthermore, quantum criticality has already been observed in such systems [295, 334, 335, 297, 336, 337]. Unlike optical schemes [338, 339, 340, 341, 342, 47], transport devices rely on current-based measurements typically operating in strongly interacting, non-Markovian regimes. This necessitates novel approaches to bound ultimate precision via the quantum Fisher information and to quantify precision of experimentally accessible measurements. Recent progress has laid this foundation [39, 343], positioning transport as a natural arena for quantum metrology—experimentally ubiquitous and supported by mature measurement infrastructure. An exciting avenue for exploration lies in the emerging connection between quantum thermodynamics [344] and metrology, where recent investigations have begun to reveal deep conceptual and practical links [345, 346, 347, 348, 303].
  • Continuous Measurements. A recurring message from recent work on continuously monitored quantum systems is that the information contained in the full measurement record is substantially richer than what can be extracted from the time-integrated or time-local signal alone. For both homodyne and photodetection schemes, Refs. [349, 350] show that the information contained in the mean signal captures only part of the available sensitivity, while the full Fisher information arises from temporal correlations in the measurement record. Refs. [351, 352] further demonstrate that standard time-local measurement strategies may fail to extract the full QFI of the emitted field. In such cases, exploiting temporal correlations—either through coherent-absorber schemes or by directly analyzing time correlations in the emission record—can recover the missing information, reveal enhanced scaling, and ultimately saturate the QFI, underscoring the central role of temporal structure in optimal continuous sensing. Finally, Ref. [47] analyzes the optimality of continuous measurements in critical quantum systems, identifying the conditions under which continuous monitoring can saturate the QFI and realize the enhanced precision scaling associated with criticality. These results point toward a promising direction for future quantum sensing architectures, where exploiting the full temporal structure of continuous measurement records could unlock the ultimate precision limits of critical quantum systems.
  • Global Sensing. Recent works [42, 52, 312] highlight that in critical quantum sensing, the achievable quantum enhancement depends sensitively on the available prior knowledge: any nonadaptive strategy will fail to exploit the critical advantage if the prior is insufficient, and the amount of prior knowledge required scales with the number of photons or particles. Adaptive measurement strategies with real-time feedback can overcome this limitation, enabling sub-shot-noise precision even with a limited number of measurements and substantial prior uncertainty. This underscores that practical implementations of global sensing in critical systems must carefully account for both the system size and the available prior information to fully harness the metrological potential offered by critical quantum sensing.
@@ Acknowledgements.
S.F. acknowledges financial support from the foundation Compagnia di San Paolo, grant vEIcolo no. 121319, and from PNRR MUR project PE0000023-NQSTI financed by the European Union – Next Generation EU. U.A. and R.D. acknowledge financial support from the Academy of Finland, grants no. 353832 and 349199. G.M. acknowledges fruitful exchanges with Scott and is grateful to Andrew Mitchell for suggesting the fermionic example in Sec XI. K.G. would like to acknowledge discussions with Linda and Lidia.